OpenStudy (anonymous):

factor 12v^2+36vp+27p^2

OpenStudy (angela210793):

3(4v^2+12vp+9p^2)

OpenStudy (anonymous):

can it be factored further?

OpenStudy (mathteacher1729):

Yes. Note that there is a perfect square hidden in \[3(4v^2+12vp+9p^2)\] \[3(4v^2+12vp+9p^2)=3(2v+3p)(2v+3p)\] So that means \[ 12v^2+36vp+27p^2 =3(4v^2+12vp+9p^2) =2(2v+3p)^2 \]

OpenStudy (anonymous):

3(2v+3p)^2??

OpenStudy (mathteacher1729):

To see if that is the answer, expand it. First compute \[(2v+3p)^2\] and then multiply that by three. :)